OpenStudy (anonymous):

Solve for x: (the quad root of)(6x-5)=(the quad root of)(x+10)

OpenStudy (anonymous):

Quad it (again).

OpenStudy (anonymous):

So would it look like this? (again) \[6^{4}(x)^{4}-5^{4}=x ^{4}+10^{4}?\]

OpenStudy (anonymous):

I have told u now 4 times altogether. 6x-5=x+10 so x is 1

OpenStudy (anonymous):

6(1)-5=(1)+10 6-5=1+10 1=11? I don't get you mean by "quad it" quad what?

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

Eve.rything

OpenStudy (anonymous):

If u quad the fourth root of 6x-5 u get 6x-5.

OpenStudy (anonymous):

If u quad the the fourth root of (x+10) u get x+10

OpenStudy (anonymous):

Therefore, for the fifth time 6x-5=x+10 so x is 1

OpenStudy (anonymous):

5x=15 x=3

OpenStudy (anonymous):

HA, HA, AND HA!

OpenStudy (anonymous):

JK

OpenStudy (anonymous):

Whatever, that's just a computation.

OpenStudy (anonymous):

None the less you DID help me find the answer, so thank you! Even though you could've just said that be quading the whole thing it makes the quad root cancel out. Thank you. :)

OpenStudy (anonymous):

I did say that, several times, with examples, in the other thread. Honest:-)

OpenStudy (anonymous):

Haha, thanks 8-)